Compute the corresponding [OH-] or [H3O+] in each of the following conditions. Describe the condition as acidic or basic.




1. [H30]=2.3 x 10 M




2. [OH-] = 7.4x10-³M




3. [H30]=8.7 x 10-M

1
Expert's answer
2021-11-11T15:14:02-0500

1) 14-2.3=11.7

2) 14-7.4=6.6

3) 14-8.7=5.3